OpenStudy (anonymous):

Solve: 5(x + 8) = 7(x − 2)

OpenStudy (anonymous):

First we distribute- 5x+40=7x-14 [Now we subtract 5x on both sides] -5x -5x ------------ 40=2x-14 [Now we add 14 on both sides] +14 +14 ----------- 54/2=2x/2 [Finally we divide by 2 on both sides to isolate the x] 27=x

OpenStudy (campbell_st):

expand both sides \[5x + 40 = 7x - 14\] add 14 to both sides \[5x + 54 = 7x\] subtract 5x from both sides \[54 = 2x\] I'll let you finish the question

OpenStudy (anonymous):